OpenStudy (anonymous):

A cone has a diameter of 12 in. and a height of 16 in. What is the volume of the cone? Use 3.14 as an approximation for straight pi and express your final answer in hundredths. @egbeach

OpenStudy (anonymous):

602.88?

OpenStudy (egbeach):

\[\Pi r ^{2^{}}\frac{ h }{ 3}\] is the formula

OpenStudy (egbeach):

radius= 1/2 diameter

OpenStudy (anonymous):

yup! :) so is my answer right?

OpenStudy (egbeach):

\[6^{2}\frac{ 16 }{ 3 }\Pi \]

OpenStudy (egbeach):

im getting a different answer.

OpenStudy (egbeach):

yes your answer is correct. haha this one you had to multiply to pi

OpenStudy (anonymous):

Thank You! ^.^

OpenStudy (egbeach):

anytime